OpenStudy (anonymous):

A person pushes a box up a ramp that is 3 m long, and 1 m high. If the box has a mass of 20 kg, and the person pushes with a force of 80 n, what is the efficiency of the ramp

OpenStudy (anonymous):



OpenStudy (anonymous):

Fk=K* Fn K=Fk/Fn

OpenStudy (anonymous):

sum the forces in y and x directions

OpenStudy (anonymous):

to sum the forces in x and y you need the angle of the ramp which is not given directly but it could be found using tan function since adj and opp are given. tan(theta)=opp/adj=1/3 theta=atan(1/3)

OpenStudy (anonymous):

i hope that clear